# Borel 1 type mappings and the respective equi-families

### v1

@@ Abstract.
We investigate classes of functions from a topological space to a metric space that are related to those of Borel class 1. Following the idea defining an equi-Baire 1 family (due to Lecomte) we define the respective equi-families of functions from the considered classes. We observe that studying of equi-families can be reduced to the exploration of a single orbit map with values in a product space. We consider the closure of equi-families with respect to the topology of pointwise convergence. Finally, we investigate functions $f\colon X\times Y\to Z$, for metric spaces $X,Y,Z$, with sections that are equi-continuous, equi-Baire 1 or equi-Borel measurable of class $\alpha$. In particular, we generalize a result of Grande.
@@ 1. Introduction
Let $f$ be a function between two metric spaces $X$ and $Y$. We say that $f$ is Baire 1 if $f$ is the limit of a pointwise convergent sequence of continuous functions $f_{n}\colon X\to Y$, $n\in\mathbb{N}$. We say that $f$ is Borel 1 if it is $F_{\sigma}$-measurable, that is, its preimage of every open set is $F_{\sigma}$. Note that every Baire 1 function is Borel 1, but the converse does not hold in general; see [Kec95, 24B]. However, if $Y=\mathbb{R}$ or $Y$ is a Banach space, then the Baire 1 functions and the Borel 1 functions coincide; see [Kec95] and also [Fos93] and [Spu10, Remark 5.6] for more general results. Note that several authors (see, e.g. [Kec95], [Lec03], [LTZ01]) use the name Baire 1 for $F_{\sigma}$-measurable functions which can sometimes lead to confusion.
Many nice characterizations of Baire 1 (or Borel 1) functions are known. Let us mention some of them. The classical theorem of Baire states that, if $X$ is complete and $Y$ is separable, then $f$ is Borel 1 if and only if $f\upharpoonright P$ has a point of continuity for every non-empty closed set $P\subseteq X$ (see [Kur66]); for some generalizations, see [Kou93]). This last condition is called the Point of Continuity Property (PCP).
Lee, Tang, and Zhao [LTZ01] obtained an interesting $\varepsilon$-$\delta$ characterization of Borel 1 functions between two Polish spaces. The respective condition characterizing Borel 1 functions was called LTZ-property in [BHH25]. The result of [LTZ01] was later generalized in [FC14] to the case when $X$ and $Y$ are separable spaces. The authors of [FC14] used an old Lebesgue theorem [Leb05] in their proof. Although the Lebesgue result was concerned with the case $X=Y=\mathbb{R}$, it remains true if $X$ is arbitrary and $Y$ is separable [BHH25, Thm 2.1]. The respective condition stated in Lebesgue’s result was referred to in [BHH25] as the Lebesgue property.
Note that Baire 1 real-valued functions and their regular subclasses have important applications in classical theory of real functions (see [Bru94]) and functional analysis [HOR91]. Independently, some other interesting classes of maps related to Baire 1 functions have been studied, also in a general setting where $X$ is a topological space. Namely, Koumoullis [Kou93] considered fragmentable maps and functions with PCP. Bouziad [Bou12] investigated weakly separated maps, which coincide with functions with LTZ-property if $X$ is a metric space. He also used cliquish and fragmentable maps, as well as those having PCP.
@@
Our work is, in some sense, a continuation of [BKS22] and [BHH25], which were inspired by the intriguing notion of equi-Baire 1 families of functions, proposed independently by Lecomte [Lec03] and Alikhani-Koopeai [AK20]. This notion plays a role in dynamical systems similar to that of equi-continuous families (see [AK22]). We observed that other types of equi-families of functions can be of interest (see [GM06], [Gra83] and [BHH25], where equi-PCP property and equi-Lebesgue property were considered). We collect such families in Section 3 and present a simple tool which shows that the study of equi-families can be reduced to the analysis of a single orbit map with values in the product space. Finally, we solve two problems from [BKS22] concerning functions of two variables with sections being equi-continuous or equi-Borel measurable of class $\alpha$, and we obtain a generalization of the result of Grande [Gra83].
@@ 2.1. Relations between various families of functions
Implications (PCP) $\Rightarrow$ (Frag) and (Frag) $\Rightarrow$ (PCP) for hereditarily Baire spaces were proved by Koumoullis in [Kou93, Theorem 2.3]. Further, Bouziad [Bou12] showed that (Frag) $\Rightarrow$ (WS). Implication (Frag) $\Rightarrow$ (Cliq) follows easily from the definitions. Actually, Lebesgue proved that (LP) $\Rightarrow$ (Borel 1) (see [Kur66, p. 375] and [BHH25]). Lebesgue proved also that (Borel 1) $\Rightarrow$ (LP) (see also [Kur66, p. 375] and [BHH25, Theorem 2.1]) for a separable space $Y$. Implication (Cliq) $\Rightarrow$ (PWD) was proved in [Dv82]. Since every countable family $\mathcal{A}$ is $\sigma$-discrete, we have (LP) $\Rightarrow$ (GLP).
@@
Finally, we show that (GLP) $\Rightarrow$ (Cliq) if $X$ is a Baire space (see also [Hol21, proof of Theorem 3.8]). Let $f\colon X\to Y$ have the generalized Lebesgue property, $\varepsilon>0$ and $\emptyset\neq U\subseteq X$ be an open set. Similarly as in the proof of “(GLP) $\Rightarrow$ (WS)” there exists a closed cover $\mathcal{A}_{\varepsilon}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$, where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. Note that $(X_{n})_{n\in\mathbb{N}}$ is a cover of $X$ consisting of closed sets. Since $U=\bigcup_{n\in\mathbb{N}}(X_{n}\cap U)$ and $U$ is a Baire space, there is a nonempty open set $V\subseteq U$ and $n_{0}\in\mathbb{N}$ such that $V\subseteq X_{n_{0}}\cap U$. Fix $x\in V$. Since $\mathcal{A}_{n_{0}}$ is discrete, there exists a nonempty open set $V^{\prime}$ such that $V^{\prime}\cap X_{n_{0}}=V^{\prime}\cap F^{x}_{n_{0}}$. Then ${\rm diam}f(V\cap V^{\prime})\leq\varepsilon$.
@@
Assume (where it is indicated in the diagram) that $X$ is hereditarily Baire. Condition (WS) $\Leftrightarrow$ (LTZ) is an obvious observation; it was explicitly stated in introduction to [Bou12].
Implications (Frag) $\Rightarrow$ (Borel 1) and (Borel 1) $\Rightarrow$ (PCP) were proved by Koumoullis [Kou93, Theorem 4.12].
From the proof of Lee, Tang and Zhao [LTZ01] it follows that (LTZ) $\Rightarrow$ (PCP) for a hereditarily Baire $X$.
Implication (Borel 1) $\Rightarrow$ (WS) for a hereditarily Baire space $X$ was proved in [KM17, Theorem 12].
Implications (GLP) $\Rightarrow$ (LTZ) and (LTZ) $\Rightarrow$ (GLP) were proved in [KM17] (see Lemma 4 and Lemma 5, respectively).
To see (GLP) $\Rightarrow$ (LP) for separable metric spaces $X$ observe that, if $X$ is metric and separable, then a $\sigma$-discrete family must be countable. (Recall that a straightforward proof of (LTZ) $\Rightarrow$ (LP) for separable $X$ was given in [FC14, Theorem 5]). ∎
@@
Implication (WS) $\Rightarrow$ (Borel 1) for a hereditarily Baire semistratifiable space $X$ was also proved by Bouziad [Bou12, Theorem 4.5] (in fact, he proved that (WS) $\Rightarrow$ (GLP)).
The equivalence (Frag) $\Leftrightarrow$ (GLP) was proved in [Kar19, Theorem B] for a hereditarily Baire perfect paracompact space $X$.
Implication (GLP) $\Rightarrow$ (LP) works also for a collectionwise normal separable space. Every paracompact space is collectionwise normal. (For the definition of a collectionwise normal space, see [Eng89, p. 305].)
@@ 2.2. Examples
Under Martin’s Axiom and the negation of the Continuum Hypothesis, there exists an uncountable set $X\subseteq\mathbb{R}$ such that each subset of $X$ is $F_{\sigma}$ in $X$; see [Buk11, Theorem 9.6]. Assume that $Y=X$ is endowed with the discrete metric, i.e. $d(x,y)=1$ if $x\neq y$ and $d(x,y)=0$ if $x=y$. Then the function $f\colon X\to Y$, $f(x)=x$, is Borel 1. Assume that $f$ has the Lebesgue property. For $\varepsilon=\tfrac{1}{2}$ we choose a closed covering $\{F_{n}:n\in\mathbb{N}\}$ of $X$ such that ${\rm diam}f(F_{n})<\tfrac{1}{2}$. Then ${\rm diam}f(F_{n})=0$ which implies that $f(F_{n})=\{y_{n}\}$ for every $n\in\mathbb{N}$. Since the sequence $(F_{n})$ covers $X$, it follows that $Y$ is countable, a contradiction. ∎
Note that, if $X$ is a nonseparable metric space, then the identity function $\operatorname{Id}\colon X\to X$ is Borel 1 and it does not have (LP); see [BHH25, Proposition 2.1].
@@ 3. Equi-families of functions
We then present solutions of two problems from [BKS22]. Finally, we generalize the result of Grande [Gra83] showing a sufficient condition for $f\colon X\times Y\to Z$ to be a Borel function of class $\alpha$, in the language of $x$-sections and $y$-sections of $f$.
@@ 3.1. Equi-families and the orbit function
The notion of an equi-Baire 1 family introduced by [Lec03] was a starting point in the previous papers [BKS22] and [BHH25]. Let $(X,\rho)$ and $(Y,d)$ be metric spaces and $\mathcal{F}\subseteq Y^{X}$. Then $\mathcal{F}$ is called equi-Baire 1 if
@@
In the same spirit, we can modify several properties considered in Section 2 for a single function $f$ from a topological space $X$ to a metric space $(Y,d)$, and we obtain the respective equi-properties for a family $\mathcal{F}\subseteq Y^{X}$. We mention that equi-Lebesgue and equi-cliquish families were considered in [BHH25].
@@
  • is equi-fragmentable, if for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$ and for all $f\in\mathcal{F}$;
  • is equi-cliquish, if for every $\varepsilon>0$ and every open set $\emptyset\neq U\subseteq X$ there exists an open set $\emptyset\neq O\subseteq U$ such that ${\rm diam}f(O)<\varepsilon$ for all $f\in\mathcal{F}$;
  • is equi-Lebesgue, if for every $\varepsilon>0$ there is a cover $(X_{n})_{n\in\mathbb{N}}$ of $X$ consisting of closed sets such that ${\rm diam}f(X_{n})\leq\varepsilon$ for all $n\in\mathbb{N}$ and all $f\in\mathcal{F}$;
  • is equi-GLP, if for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ consisting of closed subsets of $X$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$ and all $f\in\mathcal{F}$;
  • has the point of equicontinuity property (PECP), if for every closed set $\emptyset\neq F\subseteq X$ the family $\{f\restriction_{F}\colon f\in\mathcal{F}\}$ has a point of equicontinuity.
@@
the orbit function $f^{\sharp}_{\mathscr{F}}\colon X\to Y^{T}$, where $T=|\mathscr{F}|$ (see e.g. [AG01, Sec. 4]). Assume that $Z=Y^{T}$ is equipped with the supremum metric
@@
It is easy to see that the following observation is valid. This is an extended version of Proposition 1 from [Kar25].
@@
It is enough to apply Theorem 2.4, Proposition 3.3 and Remark 3.4. Note that (2) generalizes the result of [Lec03, Proposition 32]. ∎
The following proposition extends Proposition 6.4 from [BHH25].
@@
is open for each $\varepsilon>0$; see [FHT17].
We can generalize Proposition 6.3 from [BHH25] as follows.

### v2

@@ Abstract.
We investigate classes of functions from a topological space to a metric space that are related to those of Borel class 1. Following the idea defining an equi-Baire 1 family (due to Lecomte) we define the respective equi-families of functions from the considered classes. We observe that studying of equi-families can be reduced to the exploration of a single orbit map with values in a product space. We consider the closure of equi-families with respect to the topology of pointwise convergence. Finally, we investigate functions $f\colon X\times Y\to Z$, for metric spaces $X,Y,Z$, with sections that are equi-continuous, equi-Baire 1 or have equi-generalized Lebesgue property with respect to measurable sets of class $\alpha$. In particular, we generalize a result of Grande.
@@ 1. Introduction
Let $f$ be a function between two metric spaces $X$ and $Y$. We say that $f$ is Baire 1 if $f$ is the limit of a pointwise convergent sequence of continuous functions $f_{n}\colon X\to Y$, $n\in\mathbb{N}$. We say that $f$ is Borel 1 if it is $F_{\sigma}$-measurable, that is, its preimage of every open set is $F_{\sigma}$. Note that every Baire 1 function is Borel 1, but the converse does not hold in general; see [21, 24B]. However, if $Y=\mathbb{R}$ or $Y$ is a Banach space, then the Baire 1 functions and the Borel 1 functions coincide; see [21] and also [13] and [28, Remark 5.6] for more general results. Note that several authors (see, e.g. [21], [26], [27]) use the name Baire 1 for $F_{\sigma}$-measurable functions which can sometimes lead to confusion.
Many nice characterizations of Baire 1 (or Borel 1) functions are known. Let us mention some of them. The classical theorem of Baire states that, if $X$ is complete and $Y$ is separable, then $f$ is Borel 1 if and only if $f\upharpoonright P$ has a point of continuity for every non-empty closed set $P\subseteq X$ (see [24]); for some generalizations, see [23]). This last condition is called the Point of Continuity Property (PCP).
Lee, Tang, and Zhao [27] obtained an interesting $\varepsilon$-$\delta$ characterization of Borel 1 functions between two Polish spaces. The respective condition characterizing Borel 1 functions was called LTZ-property in [4]. The result of [27] was later generalized in [11] to the case when $X$ and $Y$ are separable spaces. The authors of [11] used an old Lebesgue theorem [25] in their proof. Although the Lebesgue result was concerned with the case $X=Y=\mathbb{R}$, it remains true if $X$ is arbitrary and $Y$ is separable [4, Thm 2.1]. The respective condition stated in Lebesgue’s result was referred to in [4] as the Lebesgue property.
Note that Baire 1 real-valued functions and their regular subclasses have important applications in classical theory of real functions (see [7]) and functional analysis [16]. Independently, some other interesting classes of maps related to Baire 1 functions have been studied, also in a general setting where $X$ is a topological space. Namely, Koumoullis [23] considered fragmentable maps and functions with PCP. Bouziad [6] investigated weakly separated maps, which coincide with functions with LTZ-property if $X$ is a metric space. He also used cliquish and fragmentable maps, as well as those having PCP.
@@
Our work is, in some sense, a continuation of [5] and [4], which were inspired by the intriguing notion of equi-Baire 1 families of functions, proposed independently by Lecomte [26] and Alikhani-Koopeai [2]. This notion plays a role in dynamical systems similar to that of equi-continuous families (see [3]). We observed that other types of equi-families of functions can be of interest (see [14], [15] and [4], where equi-PCP property and equi-Lebesgue property were considered). We collect such families in Section 3 and present a simple tool which shows that the study of equi-families can be reduced to the analysis of a single orbit map with values in the product space. Finally, we solve two problems from [5] concerning functions of two variables with sections being equi-continuous or equi-Borel measurable of class $\alpha$, and we obtain a generalization of the result of Grande [15].
@@ 2.1. Relations between various families of functions
Implications (PCP) $\Rightarrow$ (Frag) and (Frag) $\Rightarrow$ (PCP) for hereditarily Baire spaces were proved by Koumoullis in [23, Theorem 2.3]. Further, Bouziad [6] showed that (Frag) $\Rightarrow$ (WS). Implication (Frag) $\Rightarrow$ (Cliq) follows easily from the definitions. Actually, Lebesgue proved that (LP) $\Rightarrow$ (Borel 1) (see [24, p. 375] and [4]). Lebesgue proved also that (Borel 1) $\Rightarrow$ (LP) (see also [24, p. 375] and [4, Theorem 2.1]) for a separable space $Y$. Implication (Cliq) $\Rightarrow$ (PWD) was proved in [9]. Since every countable family $\mathcal{A}$ is $\sigma$-discrete, we have (LP) $\Rightarrow$ (GLP).
@@
Finally, we show that (GLP) $\Rightarrow$ (Cliq) if $X$ is a Baire space (see also [17, proof of Theorem 3.8]). Let $f\colon X\to Y$ have the generalized Lebesgue property, $\varepsilon>0$ and $\emptyset\neq U\subseteq X$ be an open set. Similarly as in the proof of “(GLP) $\Rightarrow$ (WS)” there exists a closed cover $\mathcal{A}_{\varepsilon}$ of $X$ such that ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}$, and $\mathcal{A}=\bigcup_{n\in\mathbb{N}}\mathcal{A}_{n}$, where $\mathcal{A}_{n}$ is discrete for each $n\in\mathbb{N}$. Note that for all $n\in\mathbb{N}$ and $x\in\bigcup\mathcal{A}_{n}$ there exists a unique $F^{n}_{x}\in\mathcal{A}_{n}$ such that $x\in F^{n}_{x}$. Let $X_{n}:=\bigcup\mathcal{A}_{n}$ for each $n\in\mathbb{N}$. Note that $(X_{n})_{n\in\mathbb{N}}$ is a cover of $X$ consisting of closed sets. Since $U=\bigcup_{n\in\mathbb{N}}(X_{n}\cap U)$ and $U$ is a Baire space, there is a nonempty open set $V\subseteq U$ and $n_{0}\in\mathbb{N}$ such that $V\subseteq X_{n_{0}}\cap U$. Fix $x\in V$. Since $\mathcal{A}_{n_{0}}$ is discrete, there exists a nonempty open set $V^{\prime}$ such that $V^{\prime}\cap X_{n_{0}}=V^{\prime}\cap F^{x}_{n_{0}}$. Then ${\rm diam}f(V\cap V^{\prime})\leq\varepsilon$.
@@
Assume (where it is indicated in the diagram) that $X$ is hereditarily Baire. Condition (WS) $\Leftrightarrow$ (LTZ) is an obvious observation; it was explicitly stated in introduction to [6].
Implications (Frag) $\Rightarrow$ (Borel 1) and (Borel 1) $\Rightarrow$ (PCP) were proved by Koumoullis [23, Theorem 4.12].
From the proof of Lee, Tang and Zhao [27] it follows that (LTZ) $\Rightarrow$ (PCP) for a hereditarily Baire $X$.
Implication (Borel 1) $\Rightarrow$ (WS) for a hereditarily Baire space $X$ was proved in [19, Theorem 12].
Implications (GLP) $\Rightarrow$ (LTZ) and (LTZ) $\Rightarrow$ (GLP) were proved in [19] (see Lemma 4 and Lemma 5, respectively).
To see (GLP) $\Rightarrow$ (LP) for separable metric spaces $X$ observe that, if $X$ is metric and separable, then a $\sigma$-discrete family must be countable. (Recall that a straightforward proof of (LTZ) $\Rightarrow$ (LP) for separable $X$ was given in [11, Theorem 5]). ∎
@@
Implication (WS) $\Rightarrow$ (Borel 1) for a hereditarily Baire semistratifiable space $X$ was also proved by Bouziad [6, Theorem 4.5] (in fact, he proved that (WS) $\Rightarrow$ (GLP)).
The equivalence (Frag) $\Leftrightarrow$ (GLP) was proved in [20, Theorem B] for a hereditarily Baire perfect paracompact space $X$.
Implication (GLP) $\Rightarrow$ (LP) works also for a collectionwise normal separable space. Every paracompact space is collectionwise normal. (For the definition of a collectionwise normal space, see [10, p. 305].)
@@ 2.2. Examples
Under Martin’s Axiom and the negation of the Continuum Hypothesis, there exists an uncountable set $X\subseteq\mathbb{R}$ such that each subset of $X$ is $F_{\sigma}$ in $X$; see [8, Theorem 9.6]. Assume that $Y=X$ is endowed with the discrete metric, i.e. $d(x,y)=1$ if $x\neq y$ and $d(x,y)=0$ if $x=y$. Then the function $f\colon X\to Y$, $f(x)=x$, is Borel 1. Assume that $f$ has the Lebesgue property. For $\varepsilon=\tfrac{1}{2}$ we choose a closed covering $\{F_{n}:n\in\mathbb{N}\}$ of $X$ such that ${\rm diam}f(F_{n})<\tfrac{1}{2}$. Then ${\rm diam}f(F_{n})=0$ which implies that $f(F_{n})=\{y_{n}\}$ for every $n\in\mathbb{N}$. Since the sequence $(F_{n})$ covers $X$, it follows that $Y$ is countable, a contradiction. ∎
Note that, if $X$ is a nonseparable metric space, then the identity function $\operatorname{Id}\colon X\to X$ is Borel 1 and it does not have (LP); see [4, Proposition 2.1].
@@ 3. Equi-families of functions
We then present solutions of two problems from [5]. Finally, we generalize the result of Grande [15] showing a sufficient condition for $f\colon X\times Y\to Z$ to be a Borel function of class $\alpha$, in the language of $x$-sections and $y$-sections of $f$.
@@ 3.1. Equi-families and the orbit function
The notion of an equi-Baire 1 family introduced by [26] was a starting point in the previous papers [5] and [4]. Let $(X,\rho)$ and $(Y,d)$ be metric spaces and $\mathcal{F}\subseteq Y^{X}$. Then $\mathcal{F}$ is called equi-Baire 1 if
@@
In the same spirit, we can modify several properties considered in Section 2 for a single function $f$ from a topological space $X$ to a metric space $(Y,d)$, and we obtain the respective equi-properties for a family $\mathcal{F}\subseteq Y^{X}$. We mention that equi-Lebesgue and equi-cliquish families were considered in [4].
@@
  • $\mathcal{F}$ is equi-fragmentable, if for every $\varepsilon>0$ and every closed set $\emptyset\neq F\subseteq X$ there exists an open set $U$ such that $U\cap F\neq\emptyset$ and ${\rm diam}f(U\cap F)<\varepsilon$ and for all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-cliquish, if for every $\varepsilon>0$ and every open set $\emptyset\neq U\subseteq X$ there exists an open set $\emptyset\neq O\subseteq U$ such that ${\rm diam}f(O)<\varepsilon$ for all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-Lebesgue, if for every $\varepsilon>0$ there is a cover $(X_{n})_{n\in\mathbb{N}}$ of $X$ consisting of closed sets such that ${\rm diam}f(X_{n})\leq\varepsilon$ for all $n\in\mathbb{N}$ and all $f\in\mathcal{F}$;
  • $\mathcal{F}$ is equi-GLP, if for every $\varepsilon>0$ there is a $\sigma$-discrete family $\mathcal{A}_{\varepsilon}$ consisting of closed subsets of $X$ such that $X=\bigcup\mathcal{A}_{\varepsilon}$ and ${\rm diam}f(A)\leq\varepsilon$ for all $A\in\mathcal{A}_{\varepsilon}$ and all $f\in\mathcal{F}$;
  • $\mathcal{F}$ has the point of equicontinuity property (PECP), if for every closed set $\emptyset\neq F\subseteq X$ the family $\{f\restriction_{F}\colon f\in\mathcal{F}\}$ has a point of equicontinuity.
@@
the orbit function $f^{\sharp}_{\mathscr{F}}\colon X\to Y^{T}$, where $T=|\mathscr{F}|$ (see e.g. [1, Sec. 4]). Assume that $Z=Y^{T}$ is equipped with the supremum metric
@@
It is easy to see that the following observation is valid. This is an extended version of Proposition 1 from [18].
@@
It is enough to apply Theorem 2.4, Proposition 3.3 and Remark 3.4. Note that (2) generalizes the result of [26, Proposition 32]. ∎
The following proposition extends Proposition 6.4 from [4].
@@
is open for each $\varepsilon>0$; see [12].
We can generalize Proposition 6.3 from [4] as follows.